the mean per capita income.is 21604 dollars per annum with a standard deviation of 727. what is the probability.that the sample.mean. wouldnsiffer from the true mean by more than 36 dollars.if a sample of 193 persons is randomly.selected
Answer #1

Solution :

Given that,

mean = \mu = 21604

standard deviation = \sigma = 727

\sigma\bar x = \sigma / \sqrt n = 727 / \sqrt 193

= 1 - P[-36 / 727 / \sqrt 193 < (\bar x - \mu \bar x) / \sigma \bar x < 36 / 727 / \sqrt 193]

= 1 - P(-0.69 < Z < 0.69)

= 1 - P(Z < 0.69) - P(Z < -0.69)

= 0.4902